11. (8 points) Let T:R + RP be given by the matrix A= [1 2 3] 0 1 2 [036] (a) Find a basis for Ker(T). (b) Find a basis for Igood hand writing plz
Answer #1

T IR² R3 be given by the matrix A = 1o 121 Lo 36 det i JE ROAD orn Tī 2 3 1 r a - ** T= A Y ม Lo36 a nta y + 3 Z7 2 y +zz I -. From © a-47 +3 z=0 => a = Z This ker (7) = {(2,Y,Z) EIR: x=Z, Y=-27) 546, -22, 2)ER: ZER} =${(1, -2, 1) : ZER? = spomf (1Again R(A) = {T(2,1,2)=(1,3, 2) ER? T(X,Y, 3) = (a +29+32, +22, 34+63 = *(1,0,0) +3(,1,3) pril+2 (3, 2, 6) span(1, 0, 0), (2,=> 4+24 =0) TEHO or 7 g =0 - te , 36 Also icot 262 = 0 => C,+2.0) -0 = q co Therefore @ I holds C, 50, q=0 - and this shows t